OpenStudy (anonymous):

a balloon rising straight up from level ground is tracked by a man with a video camera, 200 meters from the point of lift off. at the moment the video camera's angle of elevation is pi/6, the angle is increasing at the rate of 0.21 rad/min. how fast is the balloon rising at that moment (neglect the height of the camera)?..

OpenStudy (anonymous):



OpenStudy (anonymous):

First, let's work out the value of a tanθ = o/a tanPi/6 = 200/a 1/rt(3) = 200/a a = 200rt(3) \[a=200\sqrt{3}\]

OpenStudy (anonymous):

i think the length 200 meters is the length adjacent to the angle..

OpenStudy (anonymous):

Haha, good observation. I will fix it

OpenStudy (anonymous):



OpenStudy (anonymous):

its a differential calculus problem.. and im poor at differentiating...so if u can figure out this one.. your my hero... haha...

OpenStudy (anonymous):

So we have: tanθ=o/a tanθ=x/200 θ=tan^-1(x/200) Now when we differentiate this we get dθ/dx Do you get stuck at working out what to differentiate or how to actually differentiate the equations?

OpenStudy (anonymous):

both..hehe

OpenStudy (anonymous):

Haha, okay. (I will start the question again and explain what I am doing) What helps me when I do these questions is to write everything that you have as a derivative fraction. In the question you are given a rate is radians/minute. So as a fraction that is dθ/dt Therefore: \[\frac{d \theta}{dt}=0.21\] Now in the actual question we are asked to find the rate of change of the balloon's height with respect to time. As a fraction that is dx/dt (if x is taken as the distance) \[\frac{d x}{dt}=?\] Following so far?

OpenStudy (anonymous):

yes then the next is?

OpenStudy (anonymous):

can you differentiate tan^-1(x/200) for me.. as an example....

OpenStudy (anonymous):

Not quite, I will show you. So we have \[\frac{d \theta}{dt}\] and we want to find, \[\frac{d x}{dt}\] So we need to value of \[\frac{dx}{d \theta}\] to cancel out the theta on the top of the first derivative. Like this: \[\frac{dx}{dt} = \frac{d \theta}{dt} \times \frac{dx}{d \theta}\]

OpenStudy (anonymous):

I will do the differentiation to find \[\frac{dx}{d\theta}\]

OpenStudy (anonymous):

tanθ=x/200 x =200tanθ From an identity, the derivative of tan(x) is 1/(cos(x))^2 So \[\frac{dx}{d \theta}=200 \times \frac{1}{\cos^2 \theta}\] \[\frac{dx}{d \theta}= \frac{200}{\cos^2\theta}\]

OpenStudy (anonymous):

Do you know what the next step is?

OpenStudy (anonymous):

i thought the derivative of tan(x) is sec^2(x)

OpenStudy (anonymous):

The definition of the sec(x) function is 1/cos(x) Take a look at sec(A) on this site. http://en.wikipedia.org/wiki/Trigonometric_functions#Reciprocal_functions

OpenStudy (anonymous):

yea i got it now.. i dont know whats the next,. how to fiind the rate of rising...

OpenStudy (anonymous):

So, using the formula that I found ages ago: \[\frac{dx}{dt}=\frac{d\theta}{dt} \times \frac{dx}{d \theta}\] We can substitute in what we know: \[\frac{dx}{dt}=0.21 \times 200sec^2{\theta}\] And from the question, the value of theta is Pi/6 \[\frac{dx}{dt}=42sec^2{\frac{\pi}{6}}\] Think you can take it from here?

OpenStudy (anonymous):

is it possible to have an exact value without an identities

OpenStudy (anonymous):

Yes, I was wondering if you knew how to do this but I will do it for you. The only way to get exact values is to remember them so I use this online calculator to get them if I don't know them: http://www.wolframalpha.com/input/?i=sec%28Pi%2F6%29 So we car rewrite the equation \[\frac{dx}{dt}=42\frac{2}{\sqrt{3}}^2\] \[\frac{dx}{dt}=42\frac{4}{3}\] \[\frac{dx}{dt}=56\] So the rate of change of the height is 56 metres per second

OpenStudy (anonymous):

thanks....